# Inferring fine-grained migration patterns across the United States

### v1

@@ Abstract
Fine-grained migration data illuminate important demographic, environmental, and health phenomena. However, migration datasets within the United States remain lacking: publicly available Census data are neither spatially nor temporally granular, and proprietary data have higher resolution but demographic and other biases. To address these limitations, we develop a scalable iterative-proportional-fitting based method which reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data. We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 which captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data. These estimates are highly correlated with external ground-truth datasets, and improve accuracy and reduce bias relative to raw proprietary data. We publicly release MIGRATE estimates and provide a case study illustrating how they reveal granular patterns of migration in response to California wildfires.
@@ 1 Introduction
Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena. Migration patterns illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; effects of the COVID-19 pandemic [8]; and political polarization [9].
But migration data within the United States have serious limitations. The most fine-grained freely available datasets track migration at the county level. While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level. For example, research on flood-risk-induced migration uses models at the much more granular Census block level [10], arguing that this spatial granularity is necessary due to the highly localized nature of flood risk [11]. Research on neighborhoods which are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [12]. Movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 [13]. Publicly available, county-level migration datasets render these patterns entirely invisible.
@@
To address these limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates which combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data. To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm which repeatedly scales a noisy matrix to reconcile it with constraints [21], to reconcile the raw Infutor migration matrix with the more reliable Census constraints. We accommodate uncertainty in address mapping and incorporate data on births, deaths, and immigration. The output of our method is a set of yearly inferred migration matrices at the Census block group (CBG) level from 2010-2019 covering the 50 states and the District of Columbia. Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows which are currently available on the 5-year level, and 18 million times more granular than the state-state flows which are currently available on the 1-year level.
We comprehensively validate MIGRATE by comparing it to external data sources. We show that our estimates correlate well with fine-grained Census and ACS data on population counts and migration flows. We show that MIGRATE improves accuracy relative to raw Infutor data: in particular, it reduces the deviation from ground truth ACS county-to-county migration estimates by 90%. We document systematic demographic biases in the Infutor data, which overcounts rural, older, white, and home-owning populations, and show that MIGRATE almost entirely eliminates these biases. To illustrate the benefits of granular migration data, we provide a case study using our estimates to study migration in response to California wildfires, showing that MIGRATE reveals dramatic increases in out-migration which county-level data obscures. To provide a foundation for more precise migration research in the social, environmental, and health sciences, we release MIGRATE for other researchers at migrate.tech.cornell.edu.
2 Creating MIGRATE
Here we provide a high-level summary of our method for inferring fine-grained migration matrices by harmonizing Census and Infutor data. In the Appendix, we provide full details of processing the Infutor dataset (Appendix A), processing Census data (Appendix B), and harmonizing the two data sources (Appendix C).
@@ 2 Creating MIGRATE
We first preprocess the raw Infutor data into yearly migration matrices. The raw Infutor data consists of sequences of addresses for individual people. We use this information to compute the number of people moving between each pair of geographic areas. In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior. We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Appendix Table 2.
@@
Table 1: Summary statistics for raw Infutor data and Census data during the period of interest (2010-2019). We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period. Comparing the average yearly number of active individuals in the Infutor data to the Census population shows that Infutor under-counts the population. (Table 2 provides the number of active individuals in each year from 2010-2019). Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
Preprocessing Infutor requires mapping addresses to geographic areas. When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder [22, 23]. We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results. The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code. We are able to map 99.21% of all addresses in Infutor to CBGs. We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in Appendix A.
@@
Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations; then the 1-year counts of movers and non-movers by state; then the 1-year state-to-state flows; and finally the 1-year county populations. The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them: while CBG populations and state flows are estimated by the American Community Survey (ACS) and often contain large sampling errors, county populations are estimated by the Population Estimates Program (PEP) using more precise large-scale administrative datasets. We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (see Appendix D.2). To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [21, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges (Appendix C.4). The resulting migration matrices constitute our MIGRATE estimates. Appendix C provides additional methodological details on these scalings.
@@
Figure 1 depicts the MIGRATE estimates. Figure 1a plots the average out-migration rates for all CBGs in the United States in the 2010-19 period, highlighting the spatial granularity of the estimates. This granularity often reveals important patterns which are invisible at the county-county level, as we illustrate by zooming in on New York City and comparing the county-level rates obtained from ACS flows (Figure 1b) to the CBG-level out-migration rates inferred by MIGRATE (Figure 1c). In Appendix E, we include an additional analysis showing how the socioeconomic stratification of movers’ destinations changes depending on whether we look at county-level estimates or MIGRATE estimates: residents from lower-income CBGs in higher-income counties tend to move to other lower-income CBGs, but this trend is obscured at the county level.
3 Validating MIGRATE
@@
Figure 2: Validating the MIGRATE estimates. MIGRATE estimates are highly correlated with (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population). For in-migration rates, correlations are weighted by state or county population, and points are sized by population. (d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data. We compute RMSE between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates. Bars show the mean reduction in RMSE across all years; error bars plot standard deviation across years. For the in-migration rates, RMSE is weighted by area population.
@@
MIGRATE estimates are well-correlated with all available Census measures (Figure 2a-b). We first assess correlations with Census datasets incorporated in our harmonization procedure, to confirm it indeed helps reconcile MIGRATE with Census data. We would expect these correlations to be high but not uniformly perfect, because the Census datasets are not totally consistent with each other. By design, MIGRATE estimates perfectly match state and county populations (because the last step in our procedure is to match to county populations). MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations [^1] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data. We only use ACS data products whose period completely overlaps with the 2010-2019 MIGRATE range.. MIGRATE estimates are also highly correlated with state-to-state flows ($\rho=1$); this high correlation is driven in part by variation in state populations, as well as by the fact that most people do not move. Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix D.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
@@
Importantly, we further verify that our estimation procedure yields highly correlated estimates even with datasets which are not used in our estimation procedure. This constitutes a validation on held-out data, a standard check for statistical estimation methods [25]. Specifically, we compare our MIGRATE estimates to ACS 5-year county-to-county flows, which are not incorporated in our estimation, and find they are highly correlated with our estimates ($\rho=1$). This high correlation is driven in part by the fact that most people do not move; hence, in Appendix D.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$). Another contributor to the high correlation in county-county flows is variation in county size, so we separately compute correlations with county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) which normalize for county size. Correlations remain high (population-weighted $\rho=0.715$).
@@
As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation: for example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census block group populations). See Appendix D.2 for full results.
@@
Overall, these validations demonstrate that MIGRATE estimates are highly correlated with external ground-truth Census data. This is not merely because of variation in state or county size (since it remains true when examining in-migration rates, which normalize for state or county size, as well as when looking at the tract and CBG level). Nor is it due merely to the fact that most people do not move (since it remains true when examining correlations for movers specifically). Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets which are not used in our estimation procedure.
@@
MIGRATE estimates eliminate error in state and county populations because they are constructed to match county-level estimates. Relative to raw Infutor data, MIGRATE estimates reduce error in Census tract populations by an average of $85.9\%$ and error in CBG populations by $83.6\%$. MIGRATE estimates reduce error in state-to-state flows by $89.5\%$ and error in state in-migration rates by $83.3\%$. We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix D.1.
We again confirm that we still see these reductions in error on held-out data which is not used in our estimation procedure. Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$. Neither of these datasets was used in producing MIGRATE estimates. We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix D.2). Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data.
@@
Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates. Infutor data displays biases which MIGRATE estimates greatly reduce. (a) Average errors in county populations in Infutor data relative to Census data; orange denotes counties where Infutor underrepresents the population, and purple denotes counties where Infutor overrepresents it. MIGRATE estimates remove all county-level errors by construction. (b) Spearman correlation between county demographics (x-axis) and error in Infutor estimates (y-axis). Infutor’s error is correlated with racial, socioeconomic, and other demographic characteristics. (c) Comparison of demographic bias in Infutor (purple) and MIGRATE (green). Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data. MIGRATE greatly reduces biases for all demographic subgroups.
@@
Errors in the Infutor data also correlate with demographics (Figure 3b). Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups. These biases are consistent with biases found in previous research [20], as we discuss in detail in Appendix D.3, and could propagate into biases in downstream analyses of inequality and other topics.
@@
MIGRATE estimates reduce biases in Infutor data. Figure 3c compares the demographic biases of the MIGRATE estimates to the biases of the Infutor data. To quantify bias, we compare (1) the ground-truth number of people within each group (from Census data) to (2) the number of people within each group estimated from Infutor or MIGRATE, which we compute as $\sum_{i}p_{i}\cdot n_{i}$, where $i$ indexes CBGs, $p_{i}$ is the proportion of people in a CBG within a given group (e.g., the proportion of Black residents) based on Census data, and $n_{i}$ is the number of people in a CBG in Infutor or MIGRATE. This quantifies how much Infutor and MIGRATE overcount or undercount a given demographic group, relative to ground-truth Census data. The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases. We present analyses for additional demographic groups, finding similar results, in Appendix D.4.
4 Case study: migration in response to California wildfires

### v2

@@ Abstract
Fine-grained migration data illuminate important demographic, environmental, and health phenomena. However, migration datasets within the United States remain lacking: publicly available Census data are neither spatially nor temporally granular, and proprietary data have higher resolution but demographic and other biases. To address these limitations, we develop a scalable iterative-proportional-fitting based method that reconciles high-resolution but biased proprietary data with low-resolution but more reliable Census data. We apply this method to produce MIGRATE, a dataset of annual migration matrices from 2010 - 2019 that captures flows between 47.4 billion pairs of Census Block Groups — about four thousand times more granular than publicly available data. These estimates are highly correlated with external ground-truth datasets, and improve accuracy and reduce bias relative to raw proprietary data. We use MIGRATE to analyze both national and local migration patterns. Nationally, we document temporal and demographic variation in homophily, upward mobility, and moving distance: for example, we find that people are increasingly likely to move to top-income-quartile CBGs and identify racial disparities in upward mobility. We also show that MIGRATE can illuminate important local migration patterns, including out-migration in response to California wildfires, that are invisible in coarser previous datasets. We publicly release MIGRATE to provide a resource for migration research in the social, environmental, and health sciences.
@@ 1 Introduction
Fine-grained migration data, which record the number of people relocating from one geographic area to another, are essential for understanding a range of social, environmental, and health phenomena. Migration data illuminate responses to environmental disasters and climate change [1, 2, 3]; responses to economic stresses and opportunities [4, 5, 6]; patterns of social change [7]; consequences of conflicts [8]; effects of the COVID-19 pandemic [9]; and political polarization [10].
But migration datasets within the United States have serious limitations. The most fine-grained freely available datasets track migration at the county level [11]. While widely used, these datasets lack sufficient spatial resolution to study a number of phenomena where previous research has revealed important variation at the sub-county level. For example, research on flood-risk-induced migration uses models at the much more granular Census block level [12], arguing that this spatial granularity is necessary due to the highly localized nature of flood risk [13]. Research on neighborhoods that are vulnerable to climate-induced eviction or displacement similarly models risk at a sub-county (Census tract) level [14]. Movers within the same county have accounted for more than half of migratory flows in the United States in all years from 2006 to 2019 [15]. Publicly available, county-level migration datasets render these patterns entirely invisible.
@@
To address the limitations of existing migration data, we create and release MIGRATE (Migration Inference for GRAnular Trend Estimation): fine-grained migration estimates that combine the strengths of both aforementioned types of data by harmonizing biased but fine-grained proprietary data with reliable but coarse Census data. To produce our estimates, we develop a method based on iterative proportional fitting (IPF), an algorithm that repeatedly scales a noisy matrix to reconcile it with constraints [23, 24, 25], to reconcile the raw Infutor migration matrix with the more reliable Census constraints. We accommodate uncertainty in address mapping and incorporate data on births, deaths, and immigration. The output of our method is a set of yearly inferred migration matrices at the Census block group (CBG) level from 2010-2019 covering the 50 states and the District of Columbia. Our matrices capture migration flows between 47.4 billion pairs of CBGs, making MIGRATE approximately 4,600 times more granular than the county-county flows publicly available on the 5-year level, and 18 million times more granular than the state-state flows publicly available on the 1-year level.
We comprehensively validate MIGRATE by comparing it to external data sources. We show that our estimates correlate well with fine-grained Census and ACS data on population counts and migration flows. We show that MIGRATE improves accuracy relative to raw Infutor data: in particular, it reduces the deviation from ground truth ACS county-to-county migration estimates by 90%. We document systematic demographic biases in the Infutor data, which overcounts rural, older, white, and home-owning populations, and show that MIGRATE almost entirely eliminates these biases.
We then use MIGRATE to analyze both national and local migration. Nationally, we reveal both temporal and demographic variation in migration homophily, upward mobility, and moving distance. We find, for example, that people are increasingly likely to move to top-income-quartile CBGs, but also provide evidence of racial disparities: movers from plurality Black CBGs are less likely, and movers from plurality Asian CBGs more likely, to move to top-income-quartile CBGs, and these disparities persist even when controlling for income of origin CBG. Locally, we show that MIGRATE can illuminate important migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data.
Collectively, these analyses illustrate how MIGRATE reveals important migration patterns at both the national and local levels. To provide a foundation for more precise migration research in the social, environmental, and health sciences, we release MIGRATE for other researchers at migrate.tech.cornell.edu.
2 Results
2.1 Creating MIGRATE
Here we provide a high-level summary of our method for inferring fine-grained migration matrices by harmonizing Census and Infutor data. In the Methods Section, we provide full details of processing the Infutor dataset (§M.1), processing Census data (§M.2), and harmonizing both data sources (§M.3).
@@ 2.1 Creating MIGRATE
We first preprocess the raw Infutor data into yearly migration matrices. The raw Infutor data consists of sequences of addresses for individual people. We use this information to compute the number of people moving between each pair of geographic areas. In particular, we estimate the matrices $E^{(t)}$, where entries $E^{(t)}_{ij}$ represent the number of people who reside in CBG $j$ some time during year $t$ and resided in CBG $i$ a year prior [^1] We correct populations to account for birth, deaths, and international migration. Full details in the Methods Section (§M.2).. We provide descriptive statistics for the raw Infutor data and the processed migration matrices for the 2010-2019 time period in Table 1 and additional descriptive statistics in Table M1.
@@
Table 1: Summary statistics for raw Infutor data and Census data during the period of interest (2010-2019). We define “individuals with active records” as those whose [earliest Infutor date observed, latest Infutor date observed] interval intersects a given time period. Comparing the average yearly number of active individuals in the Infutor data to the Census population (second row) shows that Infutor under-counts the population. (Table M1 provides the number of active individuals in each year from 2010-2019). Active individuals have on average between two and three addresses, translating to one or two moves during the decade.
Preprocessing Infutor requires mapping addresses to geographic areas. When the raw Infutor data contains a precise street address (90.40% of all addresses in the dataset), we map each address to a unique 2010 CBG with a state-of-the-art geocoder [26, 27]. We obtain matches for 92.18% of these addresses and confirm by hand inspection of 200 addresses that this yields highly accurate results. The remaining addresses are incomplete or imprecise (e.g., contain only ZIP codes); we probabilistically map these addresses to multiple CBGs, with the weight of each CBG proportional to the population in the CBG intersecting the ZIP code. We are able to map 99.21% of all addresses in Infutor to CBGs. We fully describe the extensive additional preprocessing steps for the raw Infutor data — including accounting for uncertainty in move date, choosing between geocoders, and prioritizing addresses for individuals with conflicting histories — in the Methods Section (§M.1).
@@
Specifically, we first rescale the entries of $E^{(t)}$ to match CBG populations; then the 1-year counts of movers and non-movers by state; then the 1-year state-to-state flows; and finally the 1-year county populations. The motivation for this ordering is that the 1-year county populations are more precisely estimated than the other datasets, and we thus prioritize matching to them: while CBG populations and state flows are estimated by the American Community Survey (ACS) and often contain large sampling errors, county populations are estimated by the Population Estimates Program (PEP) using more precise large-scale administrative datasets. We verify that incorporating each of these datasets improves our performance on the validations discussed below, and that the performance is not overly sensitive to the order in which we match to the datasets (Appendix A.2). To match each matrix $E^{(t)}$ to annual county populations, we apply an IPF-based algorithm [23, 24]: we scale blocks of rows and columns to agree with the annual county populations in years $t-1$ and $t$, respectively, alternating between scaling rows and scaling columns until the procedure converges. The resulting migration matrices constitute our MIGRATE estimates. The Methods Section (§M.3) provides additional methodological details on these scalings.
@@
Figure 1 depicts the MIGRATE estimates. Figure 1a plots the average out-migration rates for all CBGs in the United States in the 2010-19 period, highlighting the spatial granularity of the estimates. This granularity often reveals important patterns which are invisible at the county-county level, as we illustrate by zooming in on New York City and comparing the county-level rates obtained from ACS flows (Figure 1b) to the CBG-level out-migration rates inferred by MIGRATE (Figure 1c).
2.2 Validating MIGRATE
@@
Figure 2: Validating the MIGRATE estimates. (a - c): MIGRATE estimates (y-axis) are highly correlated with Census data (x-axis), including (a) Census populations at the state, county, tract, and block group level, (b) Census flows between each pair of states and each pair of counties, and (c) state and county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population). For in-migration rates, correlations are weighted by state or county population, and points are sized by population. (d - f) MIGRATE estimates increase agreement with Census datasets relative to raw Infutor data. We compute RMSE between (1) MIGRATE estimates and Census data and (2) Infutor data and Census data, and report the reduction in RMSE from using MIGRATE estimates. Bars show the mean reduction in RMSE across all years; error bars plot standard deviation across years. For the in-migration rates, RMSE is weighted by area population.
@@
MIGRATE estimates are well-correlated with all available Census measures (Figure 2a-b). We first assess correlations with Census datasets incorporated in our harmonization procedure, to confirm it indeed helps reconcile MIGRATE with Census data. We would expect these correlations to be high but not uniformly perfect, because the Census datasets are not totally consistent with each other. By design, MIGRATE estimates perfectly match state and county populations (because the last step in our procedure is to match to county populations). MIGRATE estimates also achieve Pearson correlation $\rho=0.997$ with 5-year Census tract populations, and $\rho=0.996$ with 5-year CBG populations [^2] To compare our one-year MIGRATE estimates to five-year ACS estimates, we average our estimates across the same periods for consistency with the ACS data. We only use ACS data products whose period completely overlaps with the 2010-2019 MIGRATE range.. MIGRATE estimates are also highly correlated with state-to-state flows ($\rho=1$); this high correlation is driven in part by variation in state populations, as well as by the fact that most people do not move. Hence, we also compute correlation with state in-migration rates ($\rho=0.987$), which is not subject to either of these effects; in Appendix A.1, we report additional validations using diagonal and off-diagonal flow entries, as well as raw in-migration, which are also highly correlated with our estimates.
@@
Importantly, we further verify that our estimation procedure yields highly correlated estimates even with datasets which are not used in our estimation procedure. This constitutes a validation on held-out data, a standard check for statistical estimation methods [28]. Specifically, we compare our MIGRATE estimates to ACS 5-year county-to-county flows, which are not incorporated in our estimation, and find they are highly correlated with our estimates ($\rho=1$). This high correlation is driven in part by the fact that most people do not move; hence, in Appendix A.1, we compute correlations separately for county movers and county non-movers (all above $\rho=0.95$), as well as raw in-migration ($\rho=0.992$). Another contributor to the high correlation in county-county flows is variation in county size, so we separately compute correlations with county in-migration rates (i.e., the number of people moving into an area as a fraction of the area’s population) which normalize for county size. Correlations remain high (population-weighted $\rho=0.715$).
@@
As a further held-out validation, we verify that our estimation procedure yields highly correlated estimates with each Census data source even when we remove it from the datasets used for estimation: for example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data ($\rho=0.888$ with Census Tract populations and $\rho=0.856$ with Census block group populations). See Appendix A.2 for full results.
@@
Overall, these validations demonstrate that MIGRATE estimates are highly correlated with external ground-truth Census data. This is not merely because of variation in state or county size (since it remains true when examining in-migration rates, which normalize for state or county size, as well as when looking at the tract and CBG level). Nor is it due merely to the fact that most people do not move (since it remains true when examining correlations for movers specifically). Finally, it is not merely a consequence of overfitting to Census data, because it remains true on held-out datasets that are not used in our estimation procedure.
@@
MIGRATE estimates eliminate error in state and county populations because they are constructed to match county-level estimates. Relative to raw Infutor data, MIGRATE estimates reduce error in Census tract populations by an average of $85.9\%$ and error in CBG populations by $83.6\%$. MIGRATE estimates reduce error in state-to-state flows by $89.5\%$ and error in state in-migration rates by $83.3\%$. We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix A.1.
We again confirm that we still see these reductions in error on held-out data which is not used in our estimation procedure. Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$; neither of these datasets was used in producing MIGRATE estimates. We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix A.2). Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data.
@@
Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates. Infutor data displays biases that MIGRATE estimates greatly reduce. (a) Average errors in county populations in Infutor data relative to Census data; orange denotes counties where Infutor underrepresents the population, and purple denotes counties where Infutor overrepresents it. MIGRATE estimates remove all county-level errors by construction. (b) Spearman correlation between county demographics (x-axis) and error in Infutor estimates (y-axis). Infutor’s error is correlated with racial, socioeconomic, and other demographic characteristics. (c) Comparison of demographic bias in Infutor (purple) and MIGRATE (green). Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data. MIGRATE greatly reduces biases for all demographic subgroups.
@@
Errors in the Infutor data also correlate with demographics (Figure 3b). Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups. These biases are consistent with biases found in previous research [22], as we discuss in detail in Appendix A.3, and could propagate into biases in downstream analyses of inequality and other topics.
@@
MIGRATE estimates reduce biases in Infutor data. Figure 3c compares the demographic biases of the MIGRATE estimates to the biases of the Infutor data. To quantify bias, we compare (1) the ground-truth number of people within each group (from Census data) to (2) the number of people within each group estimated from Infutor or MIGRATE, which we compute as $\sum_{i}p_{i}\cdot n_{i}$, where $i$ indexes CBGs, $p_{i}$ is the proportion of people in a CBG within a given group (e.g., the proportion of Black residents) based on Census data, and $n_{i}$ is the number of people in a CBG in Infutor or MIGRATE. This quantifies how much Infutor and MIGRATE overcount or undercount a given demographic group, relative to ground-truth Census data. The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases. We present analyses for additional demographic groups, finding similar results, in Appendix A.4.
2.3 Analysis of national migration patterns
Figure 4: National migration statistics. (a) Flows between ten types of CBGs — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile. Rows correspond to the origin CBG, and columns to the destination CBG; for example, the top left entry indicates that 90% of movers from plurality white CBGs move to plurality white CBGs. The final two rows report the proportion of all movers moving to CBGs of each type, and the population share living in CBGs of each type. We report averages across all years. (b) Probability of moving to a higher-median-income CBG, conditional on income decile of origin CBG, and plurality race of origin CBG. (c) Distance moved stratified by CBG type.
We use MIGRATE to analyze national patterns of migration from 2010-2019. The spatial granularity of the data affords a unique opportunity to study demographic variation in migration patterns — for example, differences in migration between higher-income and lower-income CBGs — that may be obscured by county or state-level data, which show far less demographic variation. We hence divide CBGs into 10 (overlapping) categories — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile — and stratify the migration statistics we compute by these ten categories [^4] To determine the category of each CBG every year, we use the most recent ACS 5-year demographic estimates. For example, to classify CBGs into plurality races to study moves in the 2010-2011 period, we use the ACS 2006-2010 race and ethnicity estimates..
Figure 4a plots flows between the ten categories — for example, the proportion of movers from top-income-quartile CBGs who move to top-income-quartile CBGs. This reveals substantial homophily in migration: out-movers from all ten CBG categories are more likely than movers as a whole to move to CBGs of the same category. However, the strength of this homophily varies across categories. Movers from plurality Black, Asian, and Hispanic CBGs exhibit strong homophily: they are 5.5$\times$, 14.6$\times$, and 4.3$\times$ likelier than movers as a whole to move to CBGs with the same plurality race group (Figure S3a reports relative rates for all groups). We also observe income homophily, particularly for movers within top and bottom quartile CBGs: movers from bottom-income-quartile CBGs are nearly twice as likely as movers as a whole to move to CBGs in the bottom income quartile (34% versus 18% for movers as a whole); movers from top-income-quartile CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs (53% versus 32%). In Figure S3b, we confirm that this homophily does not occur merely because many moves are local and demographics are spatially correlated: we also observe homophily when restricting to long-distance (out-of-county) moves.
Figure 4a also shows that people are likelier to move to CBGs in higher income quartiles: 32% of moves are to top-income-quartile CBGs, while only 18% are to bottom-income-quartile CBGs. This trend becomes more pronounced over the decade we study (Figure S4d) and is only partially explained by the fact that top-income-quartile CBGs account for a larger share of the population (29%) than bottom-income-quartile CBGs (21%). There are also racial disparities: movers from plurality Asian CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs; movers from plurality Black CBGs are 2.0$\times$ likelier than movers as a whole to move to bottom-income-quartile CBGs.
Figure 4b investigates whether these racial disparities persist when controlling for origin CBG median income: in particular, plotting the share of movers moving to a higher-income CBG when controlling for the mover’s origin CBG median income decile. This reveals that the probability of moving to a higher-income CBG varies substantially by race even conditional on origin CBG income: movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs. For example, movers from fifth-income-decile, plurality Asian CBGs have a 71% chance of moving to a higher-income CBG; movers from fifth-income-decile, plurality Black CBGs have a 53% chance. Figure S4 shows that the same racial disparities occur when controlling for CBG median income percentile, as opposed to decile; or when examining the probability of moving to a top- or bottom-income-quartile CBG, as opposed to moving to a higher-income CBG. Overall, we find robust and substantial racial disparities in income of destination CBG that persist even conditional on income of origin CBG.
Finally, Figure 4c provides statistics on the distance of moves. 37% of movers move less than 5 miles; 40% from 5-50 miles; and 23% more than 50 miles. Figure 4c also highlights demographic variation in these statistics, revealing that movers from plurality white CBGs, rural CBGs, and higher-income CBGs are likelier to move long distances (more than 50 miles). Appendix B.4 reports additional migration distance statistics stratified by geographic boundary (i.e., whether movers remain within the same tract, county, or state) and shows that migration distance increases over the decade we study.
Overall, these results demonstrate that fine-grained migration data illuminates important demographic variation in homophily, upward mobility, and migration distance.
2.4 Analysis of local migration patterns